OpenStudy (anonymous):

Find the circumference of a circle having a 3 inch radius. (If you could show the work,I'd appreciate it) :}

OpenStudy (diyadiya):

Circumference = \(2 \times \pi \times radius \)

OpenStudy (diyadiya):

Here Radius = 3inch

OpenStudy (diyadiya):

Just substitute it to the formula :)

OpenStudy (anonymous):

6.28318 x 3? @Diyadiya

OpenStudy (diyadiya):

Circumference = 2 x 3.14 x 3

OpenStudy (anonymous):

Would it be 12.84? @Diyadiya

OpenStudy (diyadiya):

Can you check your calculation again?

OpenStudy (anonymous):

I think I forgot the decimal,haha. 18.84? :}

OpenStudy (diyadiya):

Right :D

OpenStudy (anonymous):

Thanks! :}

OpenStudy (diyadiya):

You're Welcome! :)